woolyfrog:

V = πr2h h = 3xr = 9 V = π(3)2(9) = 81π ≈ 254.34 inches3

woolyfrog:

where did the x come from it went from pi2h and then 3xr where did the x come from?!

woolyfrog:

@563blackghost

563blackghost:

I'm guessing your solving for the volume of a cylinder cause of the formula, but where exactly did you get this solution? And is there a diagram for this problem?

woolyfrog:

http://prntscr.com/inp4vj

563blackghost:

Okay I see.

563blackghost:

If you look in the beginning the question states `"a height that is triple the radius"`. This would indicate that we do not know what the radius is so we have it as a variable `r`. Though the `x` is unusual. It might be indicating 3x `x` being the variable for the radius. @dude what do you think?

woolyfrog:

ok lol

dude:

I agree with ya ghost :3

563blackghost:

I mean if the `x` means multiply then it would be represented as a dot.

woolyfrog:

how did he get 3?

563blackghost:

`triple` = `3`. So we are multiplying the radius by `3`.

woolyfrog:

Okay I understand

563blackghost:

awesome >.< is there anything else that you don't understand?

woolyfrog:

it says Use 3.14 for pi. Does that mean for every question that I do that says that whatever answer i get i have to divide it by pi?

woolyfrog:

grammar lol

563blackghost:

When you are told to `use 3.14 for pi` it simply means to substitute it in there for whatever formula you are using being it an area for a circle, circumference etc.

563blackghost:

They state this because sometimes they will make you keep it in terms of `pi`.

woolyfrog:

Does that mean whatever answer I get I have to divide it by pi?

563blackghost:

No. When you plug in pi as `3.14` you have already completed the stated use. The final answer is your final answer.

woolyfrog:

then why does it go 81pi to 254.34 inches3

563blackghost:

That's simply becuase they kept the terms of `pi` then plugged in `pi` as `3.14`. This resorts to them having to multiply it. If you plugged `pi` in the beginning and solved all the way through you would come to the same answer.

woolyfrog:

so that means when it says use pi all I do is plug in pi

563blackghost:

yup

woolyfrog:

how did they get 4/3 pi27= 36?

563blackghost:

Well the formula for volume of a sphere is \(\large\bf{\frac{4}{3} \pi r^{3}}\). They are told the diameter of the sphere is `6`, but to find the radius you divide it by `2` and get `3`. So the `radius = 3`. So they plug that in. \(\large\bf{\frac{4}{3} \pi (3)^{3}}\) So we simplify \(\bf{3^{3}} = 3 \times 3 \times 3 = 9 \times 3 = {\large\bf{\color{red}{27}}}\) So we get \(\large\bf{\frac{4}{3} \pi (27)}\) in which we `multiply 4 by 27` then `divide by 3` to get 36 `we keep pi in terms of pi`. \(\Large\bf{\frac{4 \times 27}{3} \pi (27) \rightarrow \frac{108}{3}\pi \rightarrow \color{red}{36 \pi}}\)

563blackghost:

whoops that `pi (27)` is a mis type it is suppose to be just `pi`.

woolyfrog:

why do we do 9 times 3?

woolyfrog:

i thought we used up all the threes

woolyfrog:

lol

563blackghost:

It's because of the powers. `REMEMBER THAT HOLDING TO A POWER IS MULTIPLYING ITSELF`. \(\large\bf{3^{3} = 3 \times 3 \times 3}\) We simply start pairing to solve. So we do `3x3` first. \(\large\bf{3 \times 3 = 9}\) \(\large\bf{3 \times 3 \times 3 \rightarrow 9 \times 3}\) We are left with one 3 so we multiply that by 9. \(\large\bf{3 \times 9 \rightarrow \color{red}{27}}\)

woolyfrog:

yea lol i just got confused cause you did 3x3x3=9

woolyfrog:

Thank you

563blackghost:

No problem! If you have any other questions and I'm free just tag meh ;)

woolyfrog:

lol ok